Case 1 – New data set: ABP = 90/60 mm Hg, HR 145 Beats/min, cardiac output = 3 l/min, CVP = 0 mm Hg, PAP = 30/10 mm Hg, PCWP = 2 mm Hg. Calculate the SV and fill in the table below and answer the question below.

Mean arterial pressure (MAP)

mmHg

Systemic vascular resistance (SVR)

dynes

Mean pulmonary artery pressure (PA mean)

mmHg

Pulmonary vascular resistance (PVR)

dynes

Stroke volume (SV)

ml

Regarding Case 1. The cardiac output is lower than normal primarily because:

afterload is increased

preload is decreased

contractility is decreased

rate is increased
Answer #1

Answer:- Given that, Case-1

ABP = 90/60 mm Hg, HR 145 Beats/min, cardiac output = 3 l/min, CVP = 0 mm Hg, PAP = 30/10 mm Hg, PCWP = 2 mm Hg,

Mean arterial pressure (MAP) = (2×diastolic+systolic) / 3

Pulmonary vascular resistance (PVR) = 80(PAP-PCWP)/cardiac output

Systemic vascular resistance (SVR) = 80(MAP-CVP)/cardiac output

Stroke volume (SV) = Heart rate * cardiac output

Mean arterial pressure (MAP) Systemic vascular resistance (SVR) Mean pulmonary artery pressure (PA mean) Pulmonary vascular resistance (PVR) Stroke volume (SV) 186.6 dynes 70 mmHg 26.6 dynes 435 mlPlease rate it.......